OpenStudy (anonymous):

simplify (3m^4/2m^2n)/(15m^3n^2/16m)

OpenStudy (anonymous):

(8 m^2)/(5 n)

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

Wolfram Mathmatica

OpenStudy (anonymous):

when you divide, multiply by the reciprocal just like with numbers.

OpenStudy (anonymous):

now cancel common factors

OpenStudy (anonymous):

give me one second

OpenStudy (anonymous):

\[\frac{3m^4}{2m^2n}\times \frac{16m}{15m^3n^2}\] i made a typo this one is correct

OpenStudy (anonymous):

cancel what you can, leave the rest

OpenStudy (anonymous):

first answer is wrong by the way. if you want to use wolfram you have to be very careful about your input

OpenStudy (anonymous):

yea i was not sure about that

OpenStudy (anonymous):

ok now i got \[48m^{5}/30m ^{4}n ^{3}\]

OpenStudy (anonymous):

is that close

OpenStudy (anonymous):

now time to cancel

OpenStudy (anonymous):

\[\frac{m^5}{m^4}=m^{5-4}=m^1=m\]

OpenStudy (anonymous):

oh wait, your denominator is wrong

OpenStudy (anonymous):

should be \[\frac{48m^5}{30m^5n^3}\]

OpenStudy (anonymous):

m squared times m cubed is m to the fifth

OpenStudy (anonymous):

so m's cancel out completely

OpenStudy (anonymous):

\[8/5n ^{3}\]

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

Hello jenni, ur problem solved or not?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

u come on twiddla now

OpenStudy (anonymous):

i dont know how to use, first tell me a little bit then i can make u sure about ur problem